wiki:Research/Instrumentation/Meetings/2019/07-24

SmartPhantom meeting: Wednesday 24th July 2019, 11:00 BST

Notes:

Present: KL, LC, JB, AK, HTL

  1. Notes and actions from last meeting:
    • HTL: Inject 90Sr electrons to estimate test-rig light yeild.
      • Done. Simulation estimating number of photo-electrons created (see slides). Now await measurement to allow comparison.
      • Agree: Make measurement of activity of source using silicon detector. This was done; the rate measured was 90kHz.
    • KL: Liaise to develop concept for polishing.
      • Ongoing. Possibly use diamond cutter after having glued two planes together to form station.
    • KL: Work within Imperial to find assembly space.
      • Done.
    • KL: Seek meeting with S.Gruber to get her input into the design.
      • Ongoing.
    • LC: Obtain chemical composition of resin and black dye.
      • Done. Carbon-black pigment. Its composition is "like" resin. KL send MuScat paper for reference.
    • LC: Develop concept for optical connector at camera end of lightguide and for potting of light guide at detector end.
      • Done. Example part has been printed. We discussed the dimensions and agreed to make it one piece.
    • KL: Liaise with FW to work out how to place aluminium-part order.
      • Done. Order may be placed on the code used for other components.
    • JB: Purchase parts for optical readout system.
      • Done. All parts for prototype are ordered. Optical system concept developed and initial sketch uploaded.
    • AK: Provide JB with project code for purchase of optical readout components.
      • Done.
  1. Mechanical design update: LC
    • See slides. Excellent discussion, next steps noted in LC's slides. Imperial my make sheet-metal stand.
    • We agreed to assemble jig at RAL because of need to measure accuracy of assembly. Implies some technical effort from RAL. LC will see if this is available.
  1. Optical readout update: JB
    • We discussed the concept that had been circulated. JB demonstrated the concept with parts (holder, lens assembly, locking ring).
    • Concept seemed good. Light tighting felt to be straightforward.
    • LC/JB: Agreed to finalise specification of interface between clear-fibre connector and optical assembly.
  1. Control system and SciWire readout update: AK
    • Will use LabView. AK has requested LabView license for himself and HTL. Work will be divided between AK and HT.
    • Need to purchase laptops to run/develop system: one for development one for development and operation of system.
    • Stepper motor with travel of 150mm will be purchased from ThorLabs. Order will include two locking rings for optical system.
      • JB will send part number to AK.
  1. Simulation update: HTL
    • Points noted above.
  1. Schedule: HTL and discussion (all) Gantt Chart
    • Discussion of schedule led to conclusion that there is a need to advance the construction etc. as far as possible.
    • Agreed to wind first prototype planes with scifi (not fishing line). Hope that these will be acceptable detectors for use.
    • We agreed to meet next Tuesday (time to be defined) to make a first baseline version of the schedule KL to suggest time.
  1. Next steps
    • Preparation of components indicates that assembly of winding jig could start after 12Aug19. Need to watch this space.
  1. DONM
    • 07Aug19
  1. AoB
    • None.

Summary of actions

  • KL: Liaise to develop concept for polishing.
  • KL send MuScat paper for reference.
  • KL: Seek meeting with S.Gruber and/or M.Puchalska to get her input into the design.
  • LC: Check re technical effort at RAL for assembly/dowling etc.
  • LC/JB: Finalise specification of interface between clear-fibre connector and optical assembly.
  • JB Provide ThorLabs part numbers of remaining optical components to AK.
  • KL: Suggest time for schedule discussion on 30Aug19.

Meeting details

Blackett 521 (The Tutorial Room) and teleconference

Phone details:

UK: 08082 380274
Austria: 0800 006612 

PIN: 5904750902

Full list of telephone contacts may be found here.

Agenda:

  1. Notes and actions from last meeting:
    • HTL: Inject 90Sr electrons to estimate test-rig light yeild. Image
      *
    • KL: Liaise to develop concept for polishing.
      *
    • KL: Work within Imperial to find assembly space.
      • Done.
    • KL: Seek meeting with S.Gruber to get her input into the design.
      • Ongoing.
    • LC: Obtain chemical composition of resin and black dye.
      *
    • LC: Develop concept for optical connector at camera end of lightguide and for potting of light guide at detector end.
      *
    • KL: Liaise with FW to work out how to place aluminium-part order.
      • Done. Order may be placed on the same code.
    • JB: Purchase parts for optical readout system. *
    • AK: Provide JB with project code for purchase of optical readout components.
      *
  1. Mechanical design update: LC
  1. Optical readout update: JB
  1. Control system and SciWire readout update: AK
  1. Simulation update: HTL
  1. Schedule: HTL and discussion (all) Gantt Chart
  1. Next steps
  1. DONM
    • 07Aug19
  1. AoB

Last modified 2 years ago Last modified on Jul 24, 2019 5:56:49 PM

Attachments (4)

Download all attachments as: .zip